# In-Context Imitation Learning via Next-Token Prediction

### v1

@@ Abstract
We explore how to enhance next-token prediction models to perform in-context imitation learning on a real robot, where the robot executes new tasks by interpreting contextual information provided during the input phase, without updating its underlying policy parameters. We propose In-Context Robot Transformer (ICRT), a causal transformer that performs autoregressive prediction on sensorimotor trajectories without relying on any linguistic data or reward function. This formulation enables flexible and training-free execution of new tasks at test time, achieved by prompting the model with sensorimotor trajectories of the new task composing of image observations, actions and states tuples, collected through human teleoperation. Experiments with a Franka Emika robot demonstrate that the ICRT can adapt to new tasks specified by prompts, even in environment configurations that differ from both the prompt and the training data. In a multi-task environment setup, ICRT significantly outperforms current state-of-the-art next-token prediction models in robotics on generalizing to unseen tasks. Code, checkpoints and data are available on https://icrt.dev.
Keywords: Multi-Task Learning, Next-Token Prediction, In-Context Learning
1 Introduction
@@
Learning-based single and multi-task robot policies have become increasingly capable [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. This improvement in robot capabilities can largely be attributed to progress in related fields, particularly in vision and language modeling. Inspired by the recent development of large language models (LLMs) and large vision models (LVMs) [11, 12, 13], which formulate natural language processing and vision problems all as next-token-prediction, recent works also have formulated robot learning as next-token-prediction problems and achieved state-of-the-art performance [7, 8, 14, 15]. Concurrently, there has been a surge in collecting large-scale robot datasets [16, 17, 18, 19, 20, 21, 22, 23] and pre-training models on these datasets [24, 25, 26, 27, 15].
@@
Despite being pre-trained on large datasets and showing some generalization ability, it is still challenging to teach these models to perform unseen tasks in different environments without additional training. New human demonstrations via teleoperation or new data collected from hand-crafted motion primitives, as well as another round of model-finetuning, are often needed to complete the new tasks. This process adds complexity to the workflow, making it challenging to apply these methods in real-world environments. Ideally, given one or a few demonstrations, the robot should be able to perform the task immediately. In their respective domains, LLMs and LVMs [11, 12, 13] have exhibited a similar ability, named in-context learning: a capability allowing the model to rapidly adapt to and recognize the task corresponding to the prompt provided at inference time without additional training.
@@
Is the in-context learning capability of next-token prediction models limited to vision and language domains? In this paper, we introduce In-Context Robot Transformer (ICRT), where we explore how next-token prediction models can be extended to perform real-world robot in-context learning. For ICRT, the context is provided as a series of robot trajectories corresponding to a new task. The model learns from this context to perform the task in a different environment configuration without requiring additional training. A robot trajectory is a sequence of image observations, robot proprioceptive states, and actions. This trajectory implicitly encodes task primitives and the objects the robot needs to interact with. The model extracts this information from the prompt and then executes actions following a similar pattern in its current environment.
Figure 1: In-Context Robot Transformer. When we want a robot to learn a new task, in many cases, either we have to program new primitives to perform the task, or we have to provide many human demonstrations to train an imitation learning model. Can a model learn the new task with few demonstrations without training? We train a next-token prediction model to perform in-context imitation learning on a real robot. In particular, the model learns from robot trajectories to perform continuous action predictions. At inference time, we prompt the model with robot sensorimotor trajectories collected by human teleoperation and provide the model with the observation of the new environment, and roll out the policy on a physical robot.
@@
Compared to existing one or few-shot imitation learning approaches, ICRT offers a simple framework that avoids complicated loss functions or key-point selection, and operates directly on robot trajectories for continuous control. Additionally, unlike existing next-token prediction models for robot learning, ICRT features a long context window, allowing it to train on multiple sensorimotor trajectories from the same task and use one or more sensorimotor trajectories as prompts during inference.
@@
Moreover, we observe that certain properties of the dataset are crucial for enabling effective in-context learning on real robots. Specifically, datasets that allow multiple tasks to be performed from the same initial observation are particularly beneficial. In such scenarios, unlike existing single-task datasets and many multi-task datasets where each environment has a unique object for the robot to interact with, the model must rely on the prompt to correctly identify the task and determine the appropriate object for interaction.
@@
    We introduce ICRT, a next-token prediction model that performs in-context learning on a real robot. ICRT takes robot’s sensorimotor trajectories on new tasks as context to perform specified tasks in unseen environment configurations.
@@
    We perform physical experiments on a Franka Emika robot at various levels of task difficulties to evaluate the in-context learning abilities of ICRT. Results suggest that ICRT can perform the unseen tasks specified by the prompt.
2 Related Works
2.1 Imitation Learning for Robotics
@@
Imitation learning is a popular and effective paradigm for equipping robots with various skills. The simplest algorithm in this domain, behavior cloning, has been successful across a wide range of tasks [28, 29, 30]. In recent years, alternative architectures such as energy-based models [31] and diffusion models [1] have also been proposed. Typically, these approaches require training a separate model for each task, although multi-task policies can be distilled from these task-specific models after training [32].
@@
Recent advancements have shown that using transformers for next-token prediction in sequence modeling has been particularly effective in both language and vision domains, especially for multi-task learning [33, 12, 34]. This approach has also been extended to robotic learning, where robot action planning is framed as a next-token prediction task using transformer-based architectures [35, 36, 7, 8, 14, 15]. In these models, observations are used to predict the next robot actions. In addition, in pursuit of developing generalist agents and robust robot policies, recent research has demonstrated that training policies on large, diverse datasets encompassing multiple tasks can lead to more robust and generalizable models [37, 38, 39, 40, 15, 5, 7, 36]. Octo [15] and OpenVLA [14] are trained on large robotic datasets, and are the state-of-the-art policies conditioning on goal images and language instructions (Octo) or just language instructions (OpenVLA). Octo employs a transformer with a diffusion head, which processes the goal conditions—language instructions or goal images—and the current image observation to predict robot actions. OpenVLA fine-tunes a pre-trained vision-language model to predict robot actions given vision observations and language instructions.
2.2 In-Context Learning
Despite the effort of training on large datasets, these policies still struggle with novel tasks or environments and often require fine-tuning. Several works have explored ways to bypass the need for model fine-tuning or to increase sample efficiency when generalizing to new tasks, leading to advances in zero-shot and few-shot imitation learning. Some approaches in meta-learning [41, 42, 43] enable few-shot imitation learning after training on a wide range of tasks, but still require fine-tuning in each new domain. Other works don’t require fine-tuning model parameters for generalizing to new tasks. Brown et al. [33] refers to this as “in-context learning” to differ from works that fine-tune the model parameters.
Many in-context learning methods often employ contrastive learning to train context encoders, which identify the most similar training tasks to the test task in the latent space [37, 44]. However, how to effectively integrate these methods within the next-token-prediction framework remains unclear. Valassakis et al. [45] achieved one-shot in-context learning by training a visual servoing network to align the robot’s end-effector with the object’s relative pose during the demonstration, but this approach requires an additional object segmentation model. Di Palo and Johns [46] introduced Keypoint Action Tokens, demonstrating in-context imitation learning using a large language model by representing demonstration trajectories as 3D coordinates with few-shot prompting. Unlike these approaches, ICRT operates without additional perception modules, processing raw image observations directly. Additionally, Vid2Robot [47] developed an encoder-decoder transformer that uses a demonstration video of a human and the current robot state as the prompt to generate robot actions. However, this method requires many auxiliary losses while ICRT uses a simple next-token prediction loss.
@@ 2.2 In-Context Learning
In this paper, we focus on enhancing next-token-prediction models to perform real-world in-context imitation learning with robots. ICRT bypasses the need for additional context encoders by directly using robot sensorimotor trajectories from new tasks as prompts for the transformer-based model. ICRT is closely related to the seminal work, One-Shot Imitation Learning [48] and Prompting Decision Transformer [49][48] predicts the next action by applying cross-attention between a demonstration sequence on a new task and the current observation, while [49] employs a short trajectory prompt to encode task-specific information for guiding policy generation in offline reinforcement learning. However, neither of these approaches considers image observations as inputs, nor do they extend beyond tasks in simulation. In contrast, ICRT does not model rewards, utilizes a significantly longer context window, and demonstrates in-context learning capabilities in physical experiments using image observations.
3 Problem Statement
@@
We investigate in-context imitation learning on a real-robot in a continuous control setting. The objective is to train a model with in-context learning capabilities using a multi-task dataset. At test time, the model can perform an unseen task in a new environment configuration by taking a few new human-teleoperated robot demonstrations as a prompt. We define environment configuration as the objects in the scene and their locations. Importantly, this is accomplished without any additional training on the new demonstrations.
@@
We define motion primitives as distinct robot motions used to complete different tasks. Each task is characterized by 1) a motion primitive and 2) the set of objects the robot interacts with using that primitive. By varying the test-time environment configuration from the one in the prompt, we evaluate the model’s ability to determine the appropriate motion primitive and identify the correct object to interact with. In this work, we consider new tasks to be tasks involving unseen objects but using motion primitives from the training data (for example, training on picking up a tiger toy and testing on picking up a cube).
@@
    The model is trained on a dataset consisting of diverse demonstrations of a single robot. Each demonstration trajectory contains observations from an RGB camera at a fixed position and a wrist-mounted RGB camera, proprioception, action, and the associated task type.
@@
    The task tested on the robot can be completed by human teleoperating the robot and is thus within the reachable workspace of the robot.
4 Approach
In this section, we first introduce the data composition to facilitate in-context imitation learning. We then introduce the transformer-based policy and its training formulation to leverage the data.
4.1 Data Formulation
For model training, we consider a dataset $\mathcal{D}$ of visuomotor trajectories $\mathcal{T}$. Each trajectory of length $t$ is a sequence of camera images $i_{t}$, proprioceptive robot states $s_{t}$, and actions $a_{t}$: $\mathcal{T}=(i_{1},s_{1},a_{1},...,i_{t},s_{t},a_{t})$. We use the absolute end-effector pose as the robot’s proprioceptive state and the delta robot end-effector pose between time steps as the action, which consists of delta translation, delta rotation and the continuous gripper action (see Appendix Section 8.4 for more detail). We assume a known grouping of the trajectories so that the dataset can be partitioned into disjoint sets of tasks $\mathcal{D}=\bigcup_{k=1}^{K}S_{k}$, with $S_{k}\cap S_{\ell}=\emptyset,\ k\neq\ell$, where $\mathcal{S}_{k}=\{\mathcal{T}_{k_{1}},...,\mathcal{T}_{k_{n}}\}$. In practice, this grouping can be retrieved from the semantic labels of the dataset. In this work, we utilize the existing large robotic dataset DROID [50] and a multi-task dataset manually collected in our robot setup, which we name ICRT-Multi-Task (ICRT-MT).
DROID [50] is a joint effort from different organizations and contains 76k real-world demonstrations. We randomly sample 10k demonstrations from DROID after filtering out demonstrations shorter than 30 steps and longer than 450 steps. DROID dataset labels the task through human-specified language instructions, which may be different for the same task. We organized the DROID data by grouping demonstrations based on their language instructions CLIP text embedding cosine similarity. Specifically, we use a threshold of 0.9 for grouping demonstrations. To further facilitate in-context learning, we make sure that each task group contains at least 4 trajectories so that there are sufficient trajectories to serve as prompts for each other. This results in roughly 2k trajectories that we use for pre-training ICRT.
Many trajectories in the DROID dataset are collected in a single-task setup, where only one task can be completed in the given environment. In such setup, the model can learn the shortcut of performing the task just conditioned on the current state and observation and never looks at the prompt, even though the prompt trajectories are similar to the current task to be performed. Therefore multi-task data is crucial for the model to learn from the prompt. We manually collected a multi-task dataset ICRT-Multi-Task (ICRT-MT) using the DROID setup (Figure. 4). This dataset has 1098 trajectories in total, and contains 29 tasks with 6 primitives: picking, pick-and-place, stacking, pushing, poking, opening and closing drawers. Objects used in the data collection and examples of the primitives are shown in Figure. 4. In ICRT-MT, each environment is set so that there exist more than 2 possible tasks for the current observation so that the model has to distinguish and learn the motion from the prompt.
During the training, for each trajectory, we independently apply vision augmentation on the image observations by augmenting the brightness and contrast. We additionally apply random crops and scaling to the side camera observation. We also apply proprioception noise sampled from a normal Gaussian distribution $\mathcal{N}(0,0.005)$. For each epoch, we randomly shuffle the order of trajectories from each task and concatenate them to form the training sequence. For each batch, we subsample for a subsequence of length $L=512$ as the input to the model, where $L$ is the sequence length defined as the number of observation, state, and action tuples. In practice, 512 steps usually contain up to 5 trajectories from the same task. We randomly select the first $k$ trajectories and label them as the prompt within the sequence. At least one complete trajectory is included in the prompt. This data grouping aims to capture inter-trajectory patterns, encouraging the model to generate action conditioned on the prompt trajectories. This approach differs from traditional behavior cloning methods, which typically use short input sequences that focus on modeling intra-trajectory behaviors.
@@ 4.1 Data Formulation
Figure 2: Method Overview: (Left) We encode the left and wrist camera observation with a pre-trained vision transformer. Additionally, we encode proprioception with a multilayer perceptron (MLP). We concatenate the visual latent and the proprioception’s latent and use attention pooling to extract a feature $f_{s}$ to represent the current state. We use another MLP to encode the action taken at the current step as the action feature $f_{a}$. (Right) We concatenate multiple trajectories of the same task and randomly sample the first k trajectories as the prompt. We encode the trajectories via a causal transformer, and the model decodes a series of tokens. We decode the tokens that are at the position of the state features to generate the next $h=16$ action via a MLP.
4.2 Model Architecture

### v2

# references.bib

@@ Abstract
We explore how to enable in-context learning capabilities of next-token prediction models for robotics, allowing the model to perform novel tasks by prompting it with human teleop demonstration examples without fine-tuning. We propose In-Context Robot Transformer (ICRT), a causal transformer that performs autoregressive prediction on sensorimotor trajectories, which include images, proprioceptive states, and actions. This approach allows flexible and training-free execution of new tasks at test time, achieved by prompting the model with demonstration trajectories of the new task. Experiments with a Franka Emika robot demonstrate that the ICRT can adapt to new tasks specified by prompts, even in environment configurations that differ from both the prompts and the training data. In a multi-task environment setup, ICRT significantly outperforms current state-of-the-art robot foundation models on generalization to unseen tasks. Code, checkpoints and data are available on https://icrt.dev.
I Introduction
@@
Learning-based single and multi-task robot policies have become increasingly capable [diffusion, zhao2023learning, lynch2023interactive, Reed2022, shah2023vint, roboagent, brohan2022rt, brohan2023rt2, chen2023palix, Driess2023]. This improvement in robot capabilities can largely be attributed to progress in related fields, particularly in vision and language modeling. Inspired by the recent development of large language models (LLMs) and large vision models (LVMs) [achiam2023gpt4, touvron2023llama, bai2023sequential], which formulate natural language processing and vision problems all as next-token-prediction, recent works also have formulated robot learning as next-token-prediction problems and achieved state-of-the-art performance [brohan2022rt, brohan2023rt2, kim2024openvlaopensourcevisionlanguageactionmodel, octo_2023]. Concurrently, there has been a surge in collecting large-scale robot datasets [depierre2018jacquard, Kalashnikov2018, Levine2018, acronym2020, shafiullah2023dobbe, fang2023rh20t, ebert2021bridge, walke2023bridgedata] and pre-training models on these datasets [nair2022, xiao2022, ma2022vip, radosavovic2022, octo_2023].
@@
Despite being pre-trained on large datasets and showing some generalization ability, it is still challenging to teach these models to perform unseen tasks in different environments without additional training. New human demonstrations via teleoperation or new data collected from hand-crafted motion primitives, as well as another round of model-finetuning, are often needed to complete the new tasks. This process adds complexity to the workflow, making it challenging to apply these methods in real-world environments. Ideally, given one or a few demonstrations, the robot should be able to perform the task immediately. In their respective domains, LLMs and LVMs [achiam2023gpt4, touvron2023llama, bai2023sequential] have exhibited a similar ability, named in-context learning: a capability allowing the model to rapidly adapt to and recognize the task corresponding to the prompt provided at inference time without additional training.
@@
Is the in-context learning capability of next-token prediction models limited to vision and language domains? In this paper, we introduce In-Context Robot Transformer (ICRT), where we explore how next-token prediction models can be extended to perform real-robot in-context learning. For ICRT, the context is provided as a series of robot trajectories corresponding to a new task. The model learns from this context to perform the task in a different environment configuration without requiring additional training. A robot trajectory is a sequence of image observations, robot proprioceptive states, and actions. This trajectory implicitly encodes task primitives and the objects the robot needs to interact with. The model extracts this information from the prompt and then executes actions following a similar pattern in its current environment.
@@
Compared to existing few-shot imitation learning approaches, ICRT offers a simple framework that avoids complicated loss functions, prior knowledge, and the need to identify key points or key frames, and operates directly on raw robot trajectories for continuous control. Additionally, unlike existing next-token prediction models for robot learning, ICRT features a long context window, allowing it to train on multiple sensorimotor trajectories from the same task and use one or more sensorimotor trajectories as prompts during inference.
@@
Importantly, we observe that certain properties of the dataset are crucial for enabling in-context learning on real robots. Specifically, datasets that allow multiple tasks to be performed from the same initial observation are particularly beneficial. Unlike existing single-task datasets or many multi-task datasets where each environment has a unique object for robot interaction, these scenarios require the model to rely on the prompt to correctly identify the task and determine the appropriate object for interaction.
@@
    We introduce ICRT, a robot foundation model that performs in-context learning on a real robot, which can effectively learn from context trajectories and perform unseen tasks without training.
@@
    Physical experiments on a Franka Emika robot demonstrate that ICRT can learn from the provided context and perform the unseen tasks specified by the prompt at various generalization levels.
II Related Works
II-A Multi-Task Imitation Learning for Robotics
@@
Imitation learning is an effective paradigm for equipping robots with various skills. The simplest algorithm in this domain, behavior cloning, has been successful across a wide range of tasks [pomerleau, argall2009survey, Levine2016]. In recent years, alternative architectures such as energy-based models [Florence2021ImplicitBC] and diffusion models [diffusion] have also been proposed. Typically, these approaches require training a separate model for each task, although multi-task policies can be distilled from these task-specific models after training [ha2023scaling].
@@
Recent advancements have shown that using transformers for next-token prediction in sequence modeling has been particularly effective in both language and vision domains, especially for multi-task learning [Brown2020, touvron2023llama, liu2023llava]. In pursuit of developing generalist agents and multi-task robot policies, robot action planning is framed as a next-token prediction task using transformer-based architectures trained on large multi-task robot datasets [Rpt2023, radosavovic2024humanoid, brohan2022rt, brohan2023rt2, kim2024openvlaopensourcevisionlanguageactionmodel, octo_2023, jang2022bc, jiang2022vima, reed2022generalist, 2024rtx, shah2023vint]. Octo [octo_2023] and OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel] represent the state-of-the-art among multi-task robotic policies. Octo [octo_2023] conditions on both goal images and language instructions, utilizing a transformer architecture with a diffusion head that fuses these inputs with current image observations to predict robot actions. OpenVLA[kim2024openvlaopensourcevisionlanguageactionmodel] conditions solely on language instructions, fine-tuning a pre-trained vision-language model to predict robot actions based on visual observations and language inputs.
II-B In-Context Learning
[…]
@@ II-B In-Context Learning
In this paper, we focus on enhancing next-token-prediction models to perform real-world in-context imitation learning with robots. ICRT bypasses the need for additional context encoders by directly using robot sensorimotor trajectories from new tasks as prompts for the transformer-based model. ICRT is closely related to the seminal work, One-Shot Imitation Learning [duan2017one] and Prompting Decision Transformer [xu2022prompting][duan2017one] predicts the next action by applying cross-attention between a demonstration sequence on a new task and the current environment’s state, while [xu2022prompting] employs a short trajectory prompt to encode task-specific information for guiding policy generation in offline reinforcement learning, using full state information and known reward functions. While both of these methods show their effectiveness in simulation, it is hard to have full-state information and known reward functions for all real-world robot manipulation tasks. To address these challenges, ICRT does not model rewards, utilizes a significantly longer context window, and demonstrates in-context learning capabilities in physical experiments using image observations.
Figure 2: Our physical setup with the Franka Emika robot, the wrist and side camera and the objects used in training and evaluation. We consider 6 primitives for training and choose “pick up and place” and “poke” as the primitives for evaluation (dark green).
III Problem Statement
@@
We consider in-context imitation learning under a real-robot manipulation setting. The goal is to train a model with in-context learning capabilities using a multi-task robotic dataset. At test time, the model can handle unseen tasks in novel environment configurations by using a few new human-teleoperated robot demonstrations as prompts. Here, environment configuration refers to the objects present in the scene and their spatial arrangement. Notably, this process is achieved without any additional training on the new demonstrations.
@@
We define motion primitives as distinct robot actions utilized to accomplish various tasks. Each task is characterized by 1) a specific motion primitive and 2) the set of objects the robot interacts with using that primitive. By altering the environment configuration at test time compared to the one in the prompt, we assess the model’s ability to select the appropriate motion primitive and identify the correct object for interaction. In this work, we consider new tasks to be tasks involving unseen objects but using motion primitives from the training data (for example, training on picking up a tiger toy and testing on picking up a cube).
@@
    The model is trained on a diverse multi-task demonstration dataset. Each trajectory contains RGB observations from a fixed camera and a wrist-mounted camera, proprioception and action.
@@
    The task tested on the robot is within the reachable workspace of the robot.
@@
Figure 3: Method Overview: (Left) We encode camera observations with a pre-trained vision transformer. Additionally, we encode proprioception with an MLP. We concatenate the visual latent and the proprioception’s latent and use attention pooling to extract a feature $f_{s}$ as the current state representation. We encode the current action with an MLP to get $f_{a}$. (Right) We concatenate multiple trajectories of the same task and randomly sample the first $k$ trajectories as the prompt. A causal transformer autoregressively predicts the next series of tokens. We decode the tokens that are at the position of the state features to generate the next $h=16$ action via an MLP.
IV Approach
In this section, we first introduce the data composition to facilitate in-context imitation learning. We then introduce the architecture and training objective for the transformer-based policy to effectively leverage the data.
IV-A Data Formulation
For model training, we consider a dataset $\mathcal{D}$ of visuomotor trajectories $\mathcal{T}$. Each trajectory of length $t$ is a sequence of camera images $i_{t}$, proprioceptive robot states $s_{t}$, and actions $a_{t}$: $\mathcal{T}=(i_{1},s_{1},a_{1},...,i_{t},s_{t},a_{t})$. We use the absolute end-effector pose as the robot’s proprioceptive state and the delta robot end-effector pose between time steps as the action, which consists of delta translation, delta rotation and the continuous gripper action. We assume a known grouping of the trajectories so that the dataset can be partitioned into disjoint sets of tasks $\mathcal{D}=\bigcup_{k=1}^{K}S_{k}$, with $S_{k}\cap S_{\ell}=\emptyset,\ k\neq\ell$, where $\mathcal{S}_{k}=\{\mathcal{T}_{k_{1}},...,\mathcal{T}_{k_{n}}\}$. In practice, this grouping can be retrieved from the semantic labels of the dataset. In this work, we utilize the existing large robotic dataset DROID [khazatsky2024droid] and a multi-task dataset manually collected in our robot setup, which we name ICRT-Multi-Task (ICRT-MT).
DROID [khazatsky2024droid] is a joint effort from different organizations and contains 76k real-world demonstrations. We randomly sample 10k demonstrations from DROID after filtering out demonstrations shorter than 30 steps and longer than 450 steps. DROID dataset labels the task through human-specified language instructions, which may be different for the same task. We organized the DROID data by grouping demonstrations based on their language instructions CLIP text embedding cosine similarity. Specifically, we use a threshold of 0.9 for grouping demonstrations. To further facilitate in-context learning, we make sure that each task group contains at least 4 trajectories so that there are sufficient trajectories to serve as prompts for each other. This results in roughly 2k trajectories that we use for pre-training ICRT.
Many trajectories in the DROID dataset are collected in a single-task setup, where only one task can be completed in a scene (e.g. only one object is presented). In such a setup, the model can learn a shortcut solution to perform the task, by only focusing on the current observation but not the prompt. Therefore, multi-task data is crucial for the model to learn from the prompt. We manually collected a multi-task dataset ICRT-Multi-Task (ICRT-MT) using the DROID setup (Figure. 2). This dataset has 1098 trajectories in total, and contains 29 tasks with 6 primitives: picking, pick-and-place, stacking, pushing, poking, opening and closing drawers. Objects used in the data collection and examples of the primitives are shown in Figure. 2. In ICRT-MT, each environment is set so that there exist more than 2 possible tasks for the current observation so that the model has to distinguish and learn the motion from the prompt.
During the training, for each trajectory, we independently apply vision augmentation on the image observations by augmenting the brightness and contrast. We additionally apply random crops and scaling to the side camera observation. We also apply proprioception noise sampled from a normal Gaussian distribution $\mathcal{N}(0,0.005)$. For each epoch, we randomly shuffle the order of trajectories from each task and concatenate them to form the training sequence. For each batch, we sample a subsequence of length $L=512$ as the input to the model, where $L$ is the sequence length defined as the number of observation, state, and action tuples. In practice, 512 steps usually contain up to 5 trajectories from the same task. We randomly select the first $k$ trajectories and label them as the prompt within the sequence. At least one complete trajectory is included in the prompt. This data grouping aims to capture inter-trajectory patterns, encouraging the model to generate action conditioned on the prompt trajectories. This approach differs from traditional behavior cloning methods, which typically use short input sequences that focus on modeling intra-trajectory behaviors.
IV-B Model Architecture